NOBLE GAS-NOBLE METAL CHEMICAL BONDING: FOURIER TRANSFORM MICROWAVE SPEC-TROSCOPY OF Kr-COINAGE METAL MONOHALIDE COMPLEXES.

J.M. THOMAS, N.R. WALKER^{a}, L.M. REYNARD^{b} and \underline{MICHAEL C.L. GERRY}, Department of Chemistry, University of British Columbia, 2036 Main Mall, Vancouver, B. C., Canada, V6T 1Z1.

We have recently reported the discovery of weak Ar-metal chemical bonding in Ar-MX complexes (M=Cu, Ag, Au; X=F, Cl, Br)^{c}. Initial indications were that the corresponding Kr complexes are more strongly bound. This paper reports the extension to new Kr-MX complexes (M=Ag, Au; X=F, Br), for which the stronger bonding is confirmed. The complexes were prepared by ablating a precursor (e.g SF_{6}, Br_{2}) entrained as less than 1% in a Kr-containing backing gas. The complexes were formed at the mouth of a pulsed nozzle, stabilised in a supersonic jet, and characterised by Fourier Transform microwave (FTMW) rotational spectroscopy. The complexes have been found to be rigid and to have short Kr-M bonds. Drastic changes in nuclear quadrupole coupling constants, indicating significant electron rearrangement on complex formation, are greater for the Kr complexes than for their Ar counterparts. Ab initio calculations indicate donation of up to 0.2 electrons from Kr to the metal; they also produce electron density contour diagrams showing significant orbital overlap between Kr and the metal.

aCurrent address: Department of Chemistry, University of Georgia, Athens, Georgia 30602-2556, U.S.A

bCurrent address: Department of Chemistry, University of Toronto, 80 St. George Street, Toronto, Ontario, M5S 3H6, Canada

cC.J. Evans and M.C.L. Gerry J. Chem. Phys. 112 1321 (2000); C.J. Evans and M.C.L. Gerry J. Chem. Phys. 112 9363 (2000); C.J. Evans, A. Lesarri and M.C.L. Gerry J. Am. Chem. Soc. 122 6100 (2000); C.J. Evans, D.S. Rubinoff and M.C.L. Gerry Phys.

Chem. Chem. Phys. 2 3943 (2000)